OpenStudy (anonymous):

Need help with 2 algebra questions MEDAL WILL BE GIVEN!

OpenStudy (anonymous):

What do you think "vertical stretch by a factor of 3" means?

OpenStudy (anonymous):

that you do what with \[y=x^3\]?

OpenStudy (anonymous):

i remember learning about it i just don't remember what my teacher said

OpenStudy (anonymous):

vertical stretch by a factor of 3 is the same as vertical shift by a factor of 3

OpenStudy (anonymous):

Oh, no its not, sorry.

OpenStudy (anonymous):

oh, then what is it?

OpenStudy (anonymous):

f(x)+a is a vertical translation

OpenStudy (anonymous):

Sketch makes the function look taller. So it would be y=x^3 ---> y=3x^3

OpenStudy (anonymous):

ohh ok yeah

OpenStudy (anonymous):

y=(x^3-4) 4 units to the right.

OpenStudy (anonymous):

So far, y=3(x^3-4)

OpenStudy (anonymous):

So, y=3(x^3-4)-3

OpenStudy (anonymous):

ok so far i understand

OpenStudy (anonymous):

I think thats it. What is the second part of the assignment.?

OpenStudy (anonymous):

Can you draw the second problem, I am not goot at word problems.

OpenStudy (anonymous):

yup that seems right, the second one is " Wind farms are a source of renewable energy found around the world. The power P (in kilowatts) generated by a wind turbine varies directly as the cube of the wind speed v (in meters per second). If a turbine generates 500kW in a 10m/s wind, how much power does it generate in a 12m/s w "

OpenStudy (anonymous):

Can you draw this for me please? I don't know the scientific units, I need to know how to convert everything to reduce the number of labels. What formulas are you given?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

it doesnt give me formulas

OpenStudy (anonymous):

im going to open this question in a separate question